OpenStudy (jokersmart):

Enter the answer. (12^5)^=?

OpenStudy (jdoe0001):

?

OpenStudy (jokersmart):

\[\left( 12^{5} \right)^{3}\]

OpenStudy (anonymous):

what is \(5\times 3\)?

OpenStudy (jokersmart):

15

OpenStudy (jdoe0001):

\(\bf (a^n)^m\implies a^{n\cdot m}\)

OpenStudy (jokersmart):

\[12^{15}\]

OpenStudy (jdoe0001):

yeap

OpenStudy (jokersmart):

Thanks to both of you!

OpenStudy (jdoe0001):

yw